# Introducing sgboost: A Practical Guide and Implementation of sparse-group boosting in R

### v1

@@ Abstract
This paper introduces the sgboost package in R, which implements sparse-group boosting for modeling high-dimensional data with natural groupings in covariates. Sparse-group boosting offers a flexible approach for both group and individual variable selection, reducing overfitting and enhancing model interpretability. The package uses regularization techniques based on the degrees of freedom of individual and group base-learners, and is designed to be used in conjunction with the mboost package. Through comparisons with existing methods and demonstration of its unique functionalities, this paper provides a practical guide on utilizing sparse-group boosting in R, accompanied by code examples to facilitate its application in various research domains. Overall, this paper serves as a valuable resource for researchers and practitioners seeking to use sparse-group boosting for efficient and interpretable high-dimensional data analysis.
@@
group selection; variable selection; R package; Ridge regression; within-group sparsity
articletype: Application Note
@@ 1 Introduction
Regularized regression is used to model high-dimensional data to reduce the risk of overfitting and to perform variable selection. In many cases, covariates have natural groupings, such as with gene data or categorical data often found in survey data. In such cases, one may want to select whole groups of variables or just individual parts (univariate regression model). ‘sgboost‘ implements the sparse-group boosting in R and other useful functions for sparse group model interpretation unique to boosting and visualization for group and individual variable selection. The package is available on CRAN [8]. A formula object defining group base-learners and individual base-learners is utilized for model fitting. Regularization is based on the degrees of freedom $(df(\cdot))$ of individual and group base-learners as defined in [10]. For an individual base-learner, the Regularization parameter $\lambda$, and for a group base-learners the regularization parameter $\lambda^{(g)}$ is defined, such that $df(\lambda)=\alpha$ and $df(\lambda^{(g)})=1-\alpha$. The package should be used with the ’mboost’ package [4]. Sparse-group boosting is an alternative method to sparse-group lasso, employing boosted Ridge regression [14]. Although there are many methods of variable selection, most focus on group selection only, e.g. [7], [5] and [16], or individual variable selection e.g. [1], [15] and [2]. However, it should be noted, that with group variable selection with overlapping groups, one could also end up with sparse-group variable selection. There are not many R packages implementing sparse-group variable selection. There is ’SGL’ [13] implementing [14] and ’sparsegl’ [6] with a faster implementation of the sparse-group lasso.
@@
The goal of this paper is to provide a practical guide including the code on how to use sparse-group boosting in R and get the most out of the method. The code is presented within this manuscript and can be found also on GitHub together with the used dataset (https://github.com/FabianObster/sgboost-introduction).
2 Sparse-group boosting workflow with sample data
@@
We first simulate the sample data and corresponding group structure with 40 equal-sized groups. Based on a linear regression model we simulate the response variable y as part of the data.frame with $n=100$ observations and $p=200$ predictor variables (each group is formed by 5 predictors).
@@ 2.1 Defining the model
Now we use the group structure to describe the sparse group boosting formula with the function create_formula(). We only need the data.frame() describing the group structure. It should contain two variables one indicating the name of the variable in the modeling data (var_name), and one indicating the group it belongs to (group_name). Additionally, we need to pass the mixing parameter alpha and the name of the outcome variable.
@@ 2.2 Fitting and tuning the model
sgboost is to be used in conjunction with the mboost package which provides many useful functions and methods which can also be used for sparse-group boosting models.
@@ 2.3 Interpreting and plotting a sparse-group boosting model
sgboost has useful functions to understand sparse-group boosting models and reflects that final model estimates of a specific variable in the dataset can be attributed to group base-learners as well as individual base-learners depending on the boosting iteration.
@@ 2.3.1 Variable importance
A good starting point for understanding a sparse-group boosting model is the variable importance. In the context of boosting, the variable importance can be defined as the relative contribution of each predictor to the overall reduction of the loss function (negative log-likelihood). get_varimp() returns the variable importance of each base-learner/predictor selected throughout the boosting process. In the case of the selection of an individual variable - call it $x_{1}$ as well as the group it belongs to -$x_{1},x_{2},...x_{p}$, both base-learners (predictors) will have an associated variable importance as defined before. This allows us to differentiate between the individual contribution of $x_{1}$ as its own variable and the contribution of the group $x_{1}$ belongs to. It is impossible to compute the aggregated variable importance of $x_{1}$ as it is unclear how much $x_{1}$ contributes to the group. However, the aggregated coefficients can be computed using get_coef(), which also returns the aggregated importance of all groups vs. all individual variables in a separate data.frame. With plot_varimp() one can visualize the variable importance as a barplot. Since group sizes can be large, the function allows to cut of the name of a predictor after max_char_length characters. One can indicate the maximum number of predictors to be printed through n_predictors or through the minimal variable importance a predictor has to have through prop. Through both parameters, the number of printed entries can be reduced. Note, that in this case, the relative importance of groups in the legend is based only on the plotted variables and not the ones removed. Adding information about the direction of effect sizes, one could add arrows behind the bars [9]. For groups, one can use the aggregated coefficients from get_coef().
@@
Figure 2: Variable importance of the sparse-group boosting model for simulated data
In this example, we see that both individual variables and groups were selected and contributed to the reduction of the loss function. The most important predictor is the first group, followed by variable 18, and then followed by group three. This is in line with what was simulated, as variable 18 has the biggest beta value, and groups one and three are full groups, meaning all variables within the groups have a non-zero beta coefficient. Groups two and four have within-group sparsity, therefore they were selected as individual variables rather than groups.

### v2

# Sparse-Group Boosting with Balanced Selection Frequencies: A Simulation-Based Approach and R Implementation

@@ Abstract
This paper introduces a novel framework for reducing variable selection bias by balancing selection frequencies of base-learners in boosting and introduces the sgboost package in R, which implements this framework combined with sparse-group boosting. The group bias reduction algorithm employs a simulation-based approach to iteratively adjust the degrees of freedom for both individual and group base-learners, ensuring balanced selection probabilities and mitigating the tendency to over-select more complex groups. The efficacy of the group balancing algorithm is demonstrated through simulations. Sparse-group boosting offers a flexible approach for both group and individual variable selection, reducing overfitting and enhancing model interpretability for modeling high-dimensional data with natural groupings in covariates. The package uses regularization techniques based on the degrees of freedom of individual and group base-learners. Through comparisons with existing methods and demonstration of its unique functionalities, this paper provides a practical guide on utilizing sparse-group boosting in R, accompanied by code examples to facilitate its application in various research domains.
@@
sparse-group boosting; variable selection bias; R package; group-balance; within-group sparsity
@@ 1 Introduction
Regularized regression is used to model high-dimensional data to reduce the risk of overfitting and to perform variable selection. In many cases, covariates have natural groupings, such as with gene data or categorical data often found in survey data. In such cases, one may want to select whole groups of variables or just individual parts. Sparse-group boosting is a powerful statistical method that extends classical boosting methods through the incorporation of structured sparsity. This structure is particularly useful in high-dimensional settings where predictors exhibit natural groupings. ‘sgboost‘ implements the sparse-group boosting in R and other useful functions for sparse group model interpretation unique to boosting, and visualization for group and individual variable selection. The package is available on CRAN [13].
Despite their utility and predictive performance, existing variable selection techniques often lack explicit mechanisms to balance sparsity within and between groups, making sgboost as an implementation of sparse-group boosting a valuable contribution to the field of variable selection.
To address a well-known limitation in boosting methods—namely, the tendency to favor larger or more flexible groups due to inherent selection bias - we introduce a group balancing algorithm. This algorithm employs a simulation-based approach to iteratively adjust the degrees of freedom assigned to each group, thereby equalizing the selection probabilities under the null hypothesis of no association. By using the shrinkage of the $RSS$ through the degrees of freedom in ridge regression, the algorithm ensures that groups of differing sizes are penalized appropriately. This adjustment mitigates over-selection of complex groups and enhances the overall interpretability and fairness of the model. The resulting group balancing can be combined with sparse-group boosting and is integrated into sgboost via the balance() function, representing a significant methodological advancement in achieving balanced variable selection across heterogeneous groups.
The increasing availability of high-dimensional datasets such as in economics, climate research, and bioinformatics, where variable selection plays a crucial role and group structures are relevant, motivates the implementation of sgboost. Many real-world datasets contain naturally predefined groups, such as geographical regions in climate modeling, questionnaire sections in survey-based research or gene data. Traditional boosting methods struggle with these structures, either selecting too many irrelevant variables or failing to capture group-level effects. By explicitly incorporating the group structure through two-level sparsity, sgboost addresses these challenges while also enhancing predictive accuracy and model interpretability.
Sparse-group boosting [15] is an alternative method to sparse-group lasso [19], employing boosted Ridge regression. Although there are many methods of variable selection, most focus on group selection only, e.g. [12], [9] and [23], or individual variable selection e.g. [1], [22] and [6]. However, it should be noted, that in some cases of group variable selection with overlapping groups, one could also end up with sparse-group variable selection. There are not many R packages implementing sparse-group variable selection methods. There is ’SGL’ [18] implementing [19], ’sparsegl’ [10] with a faster implementation of the sparse-group lasso as well as "grpreg" [4] implementing the group exponential least absolute shrinkage and selection operator (GEL) [2], the Composite minimax concave penalty (cMCP) [3] and the group bridge [9].
@@
The goal of this paper is to provide a practical guide including the code on how to use sparse-group boosting in R and get the most out of the method. The code is presented within this manuscript and can also be found also on GitHub together with the used dataset (https://github.com/FabianObster/sgboost-introduction).
While the mboost package [8] already allows for structured regularization, sgboost is specifically designed for structural sparsity and simplifies the application of sparse-group boosting through a dedicated formula constructor, enhanced visualization, and interpretability tools.
To demonstrate the relevance of sgboost, we focus on two types of datasets: simulated data designed to reflect real-world sparsity structures and real-world data where sparse-group boosting provides clear advantages. In climate economics, the willingness of farmers to take adaptive measures against climate change is affected by multiple interdependent factors such as climatic/weather patterns, market conditions, and agronomic factors. These independent variables can be sorted into natural groups (e.g., climatic variables, economic indicators, and agronomic environment), making such data ideal for sparse-group boosting. This paper explores how sgboost can be applied and performs in these settings, highlighting its practical benefits.
2 Methods
Throughout this paper, we consider a grouped dataset $X\in\mathbb{R}^{n\times p}$ with $G$ groups and each group $g$ contains $p_{g}$ variables. We also consider a generalized linear regression setting, such as explained in [21] and [11] and their extended variations for ridge regression with a similar notation as in [20]. For simplicity, we primarily illustrate the method using linear ridge regression; however, the approach can be readily extended to generalized linear models. We refer to the parameter vector as $\beta$ and to its estimate as $\widehat{\beta}$. The linear predictor with the response function $g^{-1}(\cdot)$ together form the conditional expectation of the response variable $\mathbb{E}[y|X]=\mu=g^{-1}(X\beta)$. As we can subset the design matrix, we can estimate a model using a subset of the design matrix, which we denote as $\widehat{\beta}_{V_{g}}$. Note that with this notation $\widehat{\beta}_{V_{g}}$ is not the same as the subset of $\widehat{\beta}$ using the index set ${V_{g}}$, but rather separate estimates. We will also use the notion of a Ridge hat matrix as defined in [20] using the penalty term $\lambda\geq 0$: $H^{\lambda}=X(X^{T}X)^{-1}X^{T}$ and the degrees of freedom for which we use the definition of $\text{df}(\lambda)=\text{tr}(2H^{\lambda}-{(H^{\lambda})}^{2})$.
2.1 Sparse-Group Boosting Framework
sgboost extends traditional boosting by incorporating structured sparsity, combining component-wise and group-wise selection. This approach is particularly beneficial in high-dimensional settings such as correlated independent variables. Unlike standard boosting, which selects individual variables in isolation, sgboost allows the selection of groups or individual variables. This way, entire groups of predictors can be included or excluded, improving interpretability and predictive performance. The same holds for individual variables.
We define $p+G$ candidate sets denoted as $(V_{l}\subseteq\{1,...,p\})_{l\leq p+G}$. Each candidate set describes the indices of the variables to be considered as one group. This yields $p+G$ submatrices of the design matrix $X_{V_{l}}$ only containing the columns corresponding to the index set.
  • The first $p$ are individual base-learners only containing one variable:
    $V_{l}=\{l\}\text{ for }l\leq p,$
  • and the remaining $G$ are group base-learners with group size $p_{l}$:
    $V_{l}=\{{(v_{l})}_{1},...,{(v_{l})}_{p_{l}}\}\subseteq\{1,...,p\},\text{ for }l>p:$
Through $V_{l}$, $l\geq p$, the group structure is defined using no overlapping groups. Through this bi-level structure, within-group sparsity and between-group sparsity can be balanced. By using Ridge Regression, regularization is controlled through the degrees of freedom constraint, regulating sparsity levels. The optimal base-learner is selected based on the residual sum of squares (RSS), yielding the most informative structure either via a group or individual variable at each step.
Algorithm 1 Sparse-Group $L^{2}$ Boosting Algorithm
1:  Initialize: $m\leftarrow 0$, $\widehat{\beta}^{[0]}\leftarrow\textbf{0}_{p}$, $\widehat{\mu}^{[0]}\leftarrow X\widehat{\beta}^{[0]}$
2:  while $m<M$ do
3:     $m\leftarrow m+1$
4:     for each candidate set $V_{l}$, $l\leq p+G$ do
5:        Compute residuals: $\widehat{u}^{[m-1]}\leftarrow y-\widehat{\mu}^{[m-1]}$
6:        Fit Ridge regression:
$\widehat{\overline{\beta}}^{[m]}_{V_{l}}=((X_{V_{l}})^{T}X_{V_{l}}+\lambda_{l}I_{p})^{-1}(X_{V_{l}})^{T}(\widehat{u}^{[m-1]})$
7:        Set regularization parameter $\lambda_{l}$:
$\lambda_{l}=\begin{cases}\lambda_{l}:\text{df}(\lambda_{l})=\text{tr}(2H^{\lambda}_{V_{l}}-({H^{\lambda}_{V_{l}}})^{2})=\alpha,&l\leq p\\ \lambda_{l}:\text{df}(\lambda_{l})=\text{tr}(2H^{\lambda}_{V_{l}}-({H^{\lambda}_{V_{l}}})^{2})=1-\alpha,&l>p\end{cases}$ (1)
8:     end for
9:     Select candidate set:
$l^{*}=\arg\min_{l\leq L}(\widehat{u}^{[m-1]}-X_{V_{l}}\widehat{\overline{\beta}}_{V_{l}})^{T}(\widehat{u}^{[m-1]}-X_{V_{l}}\widehat{\overline{\beta}}_{V_{l}})$
10:     for all $l\leq p+G$ do
11:        Update coefficients:
$\widehat{\beta}^{[m]}_{{V_{l}}}=\begin{cases}\widehat{\beta}^{[m-1]}+\nu\widehat{\overline{\beta}}_{V_{l^{*}}},&l=l^{*}\\ \widehat{\beta}^{[m-1]},&l\neq l^{*}\end{cases}$
12:     end for
13:     Update estimate:
$\widehat{\mu}^{[m]}=X\widehat{\beta}^{[m]}$
14:  end while
15:  Output: $\widehat{\beta}^{[M]}$
The same algorithm can be used to fit (generalized) linear models by replacing the $L^{2}$ loss function with the modified loss function $\mathcal{L}$. This yields for individual base-learners $l\leq p$
$\mathcal{L}^{[m]}_{V_{l}}=-\sum_{i=1}^{n}\ell_{i}(\widehat{\beta}^{[m-1]}+\beta_{V_{l}})+\alpha\lambda_{l}(\beta_{V_{l}})^{T}\beta_{V_{l}},$
and for group base-learners $l>p$
$\mathcal{L}^{[m]}_{V_{l}}=-\sum_{i=1}^{n}\ell_{i}(\widehat{\beta}^{[m-1]}+\beta_{V_{l}})+(1-\alpha)\lambda_{l}(\beta_{V_{l}})^{T}\beta_{V_{l}}.$
2.2 Group adjustment
Variable selection bias can occur in the presence of grouped variables, such as categorical or functional data, making the definition of the degrees of freedom $df(\lambda)=\text{tr}(2H^{\lambda}-(H^{\lambda})^{2})$ preferable [7]. However, group selection bias can still occur because of the group size. The same issue occurs in the sparse group lasso of the group bridge, which is met by using group standardization depending on the type of regularization, such as $\sqrt{p_{g}}$ [19] [3]. Many algorithms use such an adjustment, which is also referred to as outer adjustment [5]. This is to prevent an over-selection of groups with larger group sizes. Unlike traditional methods that use fixed penalties or shrinkage parameters, this algorithm dynamically adjusts selection probabilities through repeated sampling, enabling data-driven balancing. To overcome this issue, we introduce a simulation-based algorithm that balances the selection chance of one group over another by using the degrees of freedom.
Assume we have $G$ groups, described by the index sets $V_{1},...,V_{G}$, with group sizes $p_{1},...,p_{g}$. Denote the scaling vector for the degrees of freedom as $d=(d_{1},...,d_{G})$, where each group $g$ has its own value for the degrees of freedom $d_{g}$ for $g\leq G$.
Algorithm 2 Group balancing algorithm
  Initialize: Set $r=0$ and $d_{g}^{*}=d_{g}^{[1]}\equiv c$ with constant $c\in]0,1[$ for all $g\leq G$. A reasonable starting value is $c=0.5$.
2:  for $r\leq R$ do
     $r\leftarrow r+1$
4:     Simulate $K$ versions of the outcome variable $y^{(k)}$, e.g., $y^{(k)}\sim\mathcal{N}(0_{n},I_{n})$ for $k\leq K$.
     for $k\leq K$ do
6:        Fit the learning algorithm $f:X\to y^{(k)}$ with the degrees of freedom $d^{*}$ to obtain the fitted model $\widehat{f}^{(k)}$.
     end for
8:     Retrieve the activation vector $\big{(}s^{(k)}_{1},...,s^{(k)}_{G}\big{)}\in\{0,1\}^{G}$ for each $\widehat{f}^{(k)}$, indicating selected groups. If a group is selected, the value one is assigned; if not, then the value zero.
     Compute the average selection frequency vector:
$\overline{s}=\Big{(}\overline{s}_{1},...,\overline{s}_{G}\Big{)}=\Big{(}\frac{1}{K}\sum_{k=1}^{K}s^{(k)}_{1},...,\frac{1}{K}\sum_{k=1}^{K}s^{(k)}_{G}\Big{)}^{T}.$
10:     Compute the error vector:
$c^{[r]}=\big{(}\frac{1}{G},...,\frac{1}{G}\big{)}^{T}-\overline{s}.$
     for $g\leq G$ do
12:        Update:
        if $\sum_{g=1}^{G}(c_{g}^{[r]})^{2}<\sum_{g=1}^{G}(c^{*}_{g})^{2}$ then
14:
$d_{g}^{*}$ $=d_{g}^{[r-1]}$
$d_{g}^{[r]}$ $=d_{g}^{*}+\nu c^{[r]}$
        else
16:
$\nu$ $=\gamma\nu$
$d_{g}^{[r]}$ $=(1-\eta)d_{g}^{*}+\eta(d_{g}^{[r-1]}+\nu c^{[r]}).$
        end if
18:     end for
  end for
20:  Return: $d^{*}$ as the degrees of freedom scaling vector.
Note that it is sufficient to run the boosting algorithm for only one step instead of fitting the whole algorithm, to achieve the balance, as the algorithm does not depend on the actual outcome variable. This allows for a highly efficient estimation of group preference without fully fitting the model. This compensates for simulating many repetitions of the outcome variable and refitting for each sample
The general idea is to decrease the degrees of freedom for over-selected groups and increase the degrees of freedom for under-selected groups. The step size is proportional to the imbalance, meaning strongly imbalanced groups are adjusted more than slightly imbalanced groups, and is multiplied by the learning rate $\nu$, which impacts how far the update goes away from the current estimate $d^{*}$. A larger $\nu$ leads to larger corrections, hence fewer necessary iterations, but may cause oscillations or overshooting, especially in small-sample or highly collinear settings. Choosing an appropriate learning rate is therefore a trade-off between speed of convergence and stability. Also, $K$, the number of samples of the outcome variable, increases stability and should also affect the choice of the learning rate. If the algorithm overshoots and the overall imbalance increases, a convex combination between the current best estimate and the updated parameter is used, where $\eta$ is the mixing parameter. Also, the learning rate is reduced by $\gamma\in]0,1[$, e.g. 0.9 to avoid overshooting in future steps. This approach incorporates new information while preserving the information from the previous best estimate, balancing exploration and robustness. The algorithm can be stopped after a fixed number of iterations $R$ or if $\sum_{g=1}^{G}(c_{g}^{[r]})^{2}$ is smaller than some predefined value. The algorithm does not yield a unique solution, which depends strongly on the initialization of $c$. The existence of a solution that balances the selection frequencies is guaranteed by the mean-value theorem and the law of large numbers if $K\to\infty$. One can verify that $\text{df}(\lambda_{g})\to 0$ implies $\overline{s}_{g}\to 0$ and $\text{df}(\lambda_{g})\to 1$ implies $\overline{s}_{g}\to z>0$. Furthermore, the residual sum of squares is monotonous and continuous in $df(\lambda)$. Therefore, $\frac{1}{K}\sum_{k=1}^{K}s^{(k)}_{g}=P(RSS(\lambda_{g})=max_{l\leq G}(RSS(\lambda_{l}))$ is also monotonous and continuous in $\text{df}(\lambda_{g})$. A unique solution could be achieved by fixing the degrees of freedom for one base-learner and only updating the others. Another variation of the algorithm is to perform the algorithm by updating the ridge regularization parameter $\lambda$ for each group instead of the degrees of freedom. The algorithm can be used for group boosting, but also for sparse-group boosting by expanding the group index set to include also $\tilde{V}={1},...,{p},V_{1},...V_{G}$, leading to overlapping groups. In this case, one could also update the error vector to $c^{[r]}=\big{(}\frac{\alpha}{p+G},\overset{{}^{p}}{\quad...\quad},\frac{\alpha}{p+G},\frac{1-\alpha}{p+G},\overset{{}^{G}}{\quad...\quad},\frac{1-\alpha}{p+G}\big{)}^{T}-\overline{s}$, instead of mixing the degrees of freedom. In this case, $\alpha$ would have a natural interpretation, though the odds of an individual base-learner being selected over a group base-learner as $\frac{\alpha}{1-\alpha}$. This would make the choice of $\alpha$ much easier. Then, $\alpha=0$ would still correspond to group boosting, $\alpha=1$ to componentwise boosting. The case of $\alpha=0.5$, would lead to equal selection frequencies of each base-learner, regardless of the group size and type of base-learner.
3 Results
@@
We first simulate the sample data and corresponding group structure with 40 equal-sized groups to show the sparse-group boosting workflow. Based on a linear regression model we simulate the response variable y as part of the data.frame with $n=100$ observations and $p=200$ predictor variables (each group is formed by 5 predictors).
@@ 3.1 Defining the model
Now we use the group structure to describe the sparse group boosting formula with the function create_formula(). We only need the data.frame() describing the group structure. It should contain two variables, one indicating the name of the variable in the modeling data (var_name), and one indicating the group it belongs to (group_name). Additionally, we need to pass the mixing parameter alpha and the name of the outcome variable.
@@ 3.2 Fitting and tuning the model
sgboost is to be used in conjunction with the mboost package, which provides many useful functions and methods that can also be used for sparse-group boosting models.
@@ 3.3 Interpreting and plotting a sparse-group boosting model
sgboost has useful functions to understand sparse-group boosting models, and reflects that the final model estimates of a specific variable in the dataset can be attributed to group base-learners as well as individual base-learners depending on the boosting iteration.
@@ 3.3.1 Variable importance
A good starting point for understanding a sparse-group boosting model is the variable importance. In the context of boosting, the variable importance can be defined as the relative contribution of each predictor to the overall reduction of the loss function (negative log-likelihood). get_varimp() returns the variable importance of each base-learner/predictor selected throughout the boosting process. In the case of the selection of an individual variable - call it $x_{1}$ as well as the group it belongs to -$x_{1},x_{2},...x_{p}$, both base-learners (predictors) will have an associated variable importance as defined before. This allows us to differentiate between the individual contribution of $x_{1}$ as its own variable and the contribution of the group $x_{1}$ belongs to. It is impossible to compute the aggregated variable importance of $x_{1}$ as it is unclear how much $x_{1}$ contributes to the group. However, the aggregated coefficients can be computed using get_coef(), which also returns the aggregated importance of all groups vs. all individual variables in a separate data.frame. With plot_varimp() one can visualize the variable importance as a barplot. Since group sizes can be large, the function allows for cutting of the name of a predictor after max_char_length characters. One can indicate the maximum number of predictors to be printed through n_predictors or through the minimal variable importance a predictor has to have through prop. Through both parameters, the number of printed entries can be reduced. Note, that in this case, the relative importance of groups in the legend is based only on the plotted variables and not the ones removed. Adding information about the direction of effect sizes, one could add arrows behind the bars [14]. For groups, one can use the aggregated coefficients from get_coef().
@@
Figure 2: Variable importance of the sparse-group boosting model for simulated data. The variable labels in the groups are cut off after 15 characters by defalut.
In this example, we see that both individual variables and groups were selected and contributed to the reduction of the loss function. The most important predictor is the first group, followed by variable 18, and then by group three. This is in line with what was simulated, as variable 18 has the biggest beta value, and groups one and three are full groups, meaning all variables within the groups have a non-zero beta coefficient. Groups two and four have within-group sparsity, therefore, they were selected as individual variables rather than groups.